The position of a particular particle as a function of time is given by r = (9.80t·i-885j-1.00 t2·k)m, where t is in seconds. 


 Part A

What is the average velocity of the particle between t=1.00 s and t=3.00 S? 


Part B 

What is the magnitude of the instantaneous velocity at 3.00 s?  
Answer #1

a)

at t =1

r1 = 9.6 i +8.85 j - k

at t =3

r2 = 28.8 i + 8.85 j - 9 k

displacement,

r = r2 - r1 = 19.2 i - 8 k

av velocity, = displacement / time = (19.2 i - 8k)/(3-1)

v = 9.6 i - 4 k

===========

b)

v = dr / dt

v = 9.6 i - 2t k

at t = 3

v = 9.6 i - 6 k

magnitude,

v = sqrt ( 9.6^2 + 6^2)

v = 11.321 m/s

=========

do comment in case any doubt, will reply for sure.. Goodluck